PROTON DONOR/ACCEPTOR PROPENSITIES OF AMMONIA: ROTATIONAL STUDIES OF ITS MOLECULAR COMPLEXES WITH ORGANIC MOLECULES

BARBARA M. GIULIANO, ASSIMO MARIS, SONIA MELANDRI, LAURA B. FAVERO, LUCA EVANGELISTI and \underline{WALTHER CAMINATI}, Dipartimento di Chimica ”G. Ciamician” dell’Università, Via Selmi 2, I-40126 Bologna, Italy.

We studied the rotational spectra of the adducts of ammonia with several organic molecules, namely tert-butanol,^{a} glycidol,^{b} ethyl alcohol, anisol and 1,4-difluorobenzene. The adducts with glycidol and ethanol have been observed for both conformers of the substrate molecule. Based on the rotational and ^{14}N quadrupole coupling constants of the various complexes, we found a considerably different behaviour of ammonia, with respect to water, in its proton donor/acceptor double role. In the interaction with the three alcohol molecules, NH_{3} acts as a proton acceptor and the OH groups as a proton donor. However, in the case of glycidol-NH_{3}, a secundary N-H⋯O interaction occurrs between ammonia and the ether oxygen. This interaction generates a sizable V_{3} barrier to the internal rotation of the NH_{3} moiety, while NH_{3} undergoes a free rotation in tert-butanol-NH_{3} and in ethanol-NH_{3}. As to the anisole-NH_{3} and 1,4-difluorobenzene-NH_{3} complexes, the NH_{3} group explicits its double proton donor/acceptor role, although through two weak (C_{Me}-H⋯N and N-H⋯ π) H-bonds. There is, however, an important difference between the two complexes, because in the first one NH_{3} lies out of the aromatic plane, while in the second one it is in the plane of the aromatic ring.

aB. M. Giuliano, M. C. Castrovilli, A. Maris, S. Melandri, W. Caminati and E. A. Cohen, Chem.Phys.Lett., 2008, 463, 330

bB. M. Giuliano, S. Melandri, A. Maris, L. B. Favero and W. Caminati, Angew.Chem.Int.Ed., 2009, 48, 1102